# Efficient Composite Infrared Spectroscopy: Combining the Doubly-Harmonic Approximation with Machine Learning Potentials

### v1

@@ 3.4 Performance for simulation of gas-phase IR spectra
For the IR7193 set, we investigated all combinations of the discussed levels of theory. Two sets of reference data were employed: unscaled B3LYP-3c and the respective mass-scaled results, which closely correspond to experimental data.20 For brevity, the following discussion focuses on the $r_{\mathrm{msc}}$ matchscore metric, with other metrics provided in the Supporting Information. Results are summarized in Figures 8 to 10, illustrating the average $r_{\mathrm{msc}}$ over IR7193 for all method combinations, referencing different frequency scaling for either reference or actual data.
@@
Figure 8 presents the average $r_{\mathrm{msc}}$ with respect to unscaled B3LYP-3c reference spectra. As expected, there is a correlation between previously observed performance for frequencies, dipole moment prediction, and the similarity scores for IR spectra. All MACE-OFF23 variants consistently outperform the GFNn-xTB methods, regardless of the origin of the dipole data employed. The best average $r_{\mathrm{msc}}$, with a value of 0.813, is provided by MACE-OFF23(large) in combination with MACE-$\mu$ based IR intensities. Combinations of MACE-OFF23 with dipole derivatives obtained from either of the xTB methods also show high $r_{\mathrm{msc}}$ values, around 0.79. Unfortunately, IR intensities obtained via the classical EEQ model show much poorer performance, reflecting the previously observed lack of accuracy in dipole moment prediction. The semi-classical CEH model performs between the xTB and EEQ-derived IR predictions, contrary to the similar behavior of this method for dipole moments compared to GFN1-xTB (see section 3.3). In general, similar trends are observed for the combination of GFNn-xTB frequencies with various dipole moment methods: ML and xTB-derived IR intensities are accurate, while classical and CEH-based IR intensities fail to achieve adequate accuracy. Employing a force-field, even one trained for frequency prediction such as GFN-FF, notably worsens the obtained IR spectra. Multiple factors probably contribute to this performance. Most importantly, GFN-FF shows a significant number of large RMSD minima after optimisation, as discussed in section 3.1, which naturally affects the predicted frequencies and dipoles and hence must contribute to the poor results. The best predictions are again obtained in combination with GFNn-xTB or MACE-$\mu$ calculated IR intensities, although their computational costs exceed those of calculating the GFN-FF Hessian.
Employing mass-scaled reference data to better reproduce experimental gas-phase spectra20 has a significant impact on the observed performance, as shown in Figure 9.
@@
All MACE-OFF23-based calculations show substantially poorer $r_{\mathrm{msc}}$ values, regardless of the employed dipole moment methods, which can be clearly assigned to the shift in frequencies. On the other hand, all GFN methods are much less affected by the changing reference. GFN2-xTB performs better than its predecessor GFN1-xTB, and even has slightly higher $r_{\mathrm{msc}}$ than with the plain B3LYP-3c reference. Apparently, the methods of the GFN method family are more consistent concerning the molecular frequencies (as observed already for the ZPVE, see section 3.2), and hence preform similarly in both cases. This result shows that the GFN methods are a reasonable choice for inexpensively calculating frequencies since, on average, a robust performance can be expected. Unfortunately, this observation also implies a general inability of the GFN methods to achieve much better IR spectra predictions, as well as potential short-comings of the comparison metric which is discussed below.
In Figure 10 both the reference data and the low-cost simulations target IR spectra closely corresponding to the experiment.20
@@
Compared to Figure 9, the behaviour of MACE-OFF23-based predictions with regards to the mass-scaled B3LYP-3c reference can be rectified by employing the linear frequency scaling factors obtained in section 3.2. In combination with either the xTB-based dipole derivatives or MACE-$\mu$, all MACE-OFF23 MLPs consistently achieve average $r_{\mathrm{msc}}$ over 0.8, with the overall best benchmark of 0.859 result given by MACE-OFF23(large)+MACE-$\mu$. A similar trend is seen for GFN1-xTB and GFN2-xTB, with the latter theory performing better. The best average $r_{\mathrm{msc}}$ among these SQM methods is 0.787, which is achieved with GFN2-xTB in combination with MACE-$\mu$. As before, GFN-FF-based predictions of IR spectra, as well as all IR spectra calculated with underlying EEQ or CEH dipole moment derivatives are lacking accuracy. In fact, GFN-FF results scaled by the linear factor determined via the ZPVE are on average worse than both seen in Figure 8 and Figure 9, which is further testimony to an asystematic behaviour of the method. For the underlying dipole moment derivatives, classical charge-equilibrium models show bad average $r_{\mathrm{msc}}$ values throughout, independent of both reference values and methods used for frequency calculation. Hence, their use in IR spectra computation should generally be avoided. The semiclassical CEH model on the other hand, while certainly worse than either a SQM or specially trained MLP, shows promising performance. Results for this method could certainly be improved with a specialized parametrization, and could motivate further development of similar single-diagonalization SQM Hamiltonians.
@@
Unfortunately, these results can also be interpreted as a lack of precision in matchscore-based metrics like $r_{\mathrm{msc}}$. While clearly defined or linearly shifted spectra are sufficiently well compared by matchscore metrics, for asystematically behaving methods the comparison is too insensitive to accurately distinguish spectra, and will produce some average $r_{\mathrm{msc}}$ value (as seen for GFN1-xTB and GFN2-xTB, as well as their combination with MACE-$\mu$ in Figure 9). A better metric should include clear assignment of vibrational modes between reference and calculated spectra, for which, to the best of our knowledge, no automated procedure exists, although attempts exist to project the IR intensity along localized modes and use an appropriate assignment accoringly.65 Furthermore, such a metric would only work if it is known a priori that the same conformation of the same molecule is to be compared and information about the vibrational modes is accessible for the reference spectra. If the goal is, for example, the identification of unknown compounds from experimental measurements, only matchscore metrics seem suitable. Hence, due to lack of better alternatives, we continue to refer to $r_{\mathrm{msc}}$, but advise caution to the reader. On the other hand, obtaining generally mediocre values of $r_{\mathrm{msc}}$ as seen for GFNn-xTB, may signify suitability in applications where an exact order of the frequencies and modes is not overly important, for example the calculation of the ZPVE or vibrational entropy. This fitness is in line with the original GFNn-xTB design purpose12 and hence seems to be confirmed by our results. For better visual clarity about $r_{\mathrm{msc}}$, examples of a “good”, “average” and “mediocre” matchscore are given in Figure 11, depicted as an overlay between the mass-scaled B3LYP-3c reference and the MACE-OFF23(large)+MACE-$\mu$ computed IR spectra with applied ZPVE scaling factor.
@@
In general, matchscores of 0.9 or higher exhibit a close match between the two compared spectra, with only minor differences in relative intensities. Scores between 0.8 and 0.9 likewise denote good correspondence of spectra, although some frequency shifted signals may appear, as seen for modes around 3000 to 3500 cm-1 in Figure.11b. Spectra with $r_{\mathrm{msc}}$ below 0.8 commonly disagree with regard to at least one signal, which is the case for the OH vibrational mode at around 3500 cm-1 in Figure 11c, where both the frequency and intensity are overestimated by the low-cost computation. A mismatch of this type is common cause for “mediocre” or worse matchscore results, despite other regions of the spectrum (e.g. the fingerprint regime up to $\approx$1500 cm-1) showing good agreement. Spectra with overall fewer IR signals are more prone to this effect, while signals in cluttered spectra will often overlap regardless and hence contribute to $r_{\mathrm{msc}}$. For benchmark purposes with theory-to-theory comparisons being made, a new metric depending on identification and matching of the actual vibrational mode would clearly be preferable.
@@ 4 Conclusion
In this work, we tested combinations of low-cost semiempirical, force-field, and MLP methods for calculating gas-phase IR spectra within the doubly-harmonic approximation. Our approach involves benchmarking the GFN1-xTB, GFN2-xTB, GFN-FF and MACE-OFF23 models for calculating molecular geometries and associated harmonic frequencies, as well as the GFN1-xTB, GFN2-xTB, CEH, EEQ and MACE-$\mu$ models for molecular dipole moments. Our analysis centers mostly on the IR7193 set, collecting some 7193 organic molecules and their IR spectra. Linear frequency scaling factors were determined based on the ZPVE fraction relative to a specially (mass-)scaled theoretical reference, B3LYP-3c. IR spectra were obtained by combining frequency predictions with IR intensities calculated via numerically derived dipole moment derivatives.
@@
A key observation is that the MACE-OFF23 MLPs are found to be particularly effective for drug-like molecules, providing highly accurate geometries, frequencies, and IR spectra that outperform the GFNn-xTB methods. Interestingly, MACE-OFF23 appears to inherit characteristics similar to DFT methods, as evidenced by the comparable importance and magnitude of harmonic frequency scaling factors. This result is in stark contrast to the evaluated semiempirical methods, which are mostly asystematic with regard to such treatments and on average are only little affected by a linear scaling of the harmonic frequencies. In general, although the GFNn-xTB methods offer decent performance and speed for smaller system sizes and inherently include dipole moments, they often do not match the accuracy of MACE-OFF23. This result is clearly seen from the best observed average $r_{\mathrm{msc}}$ of 0.787 of GFN2-xTB versus a $r_{\mathrm{msc}}$ of 0.859 for MACE-OFF23(large), both in combination with MACE-$\mu$-based IR intensities. Both MACE-$\mu$ and GFNn-xTB exhibit decent performance for calculation of the underlying dipole moment derivatives, with the SQM methods having a computational cost advantage for drug-sized molecules. The use of classical EEQ models for this purpose should be avoided. While CEH does not reach the accuracy of SQM or MLP dipole predictions, the observed performance is a clear motivator for future research.
@@
In summary, we see an opportunity for both MLP and SQM-based (or the combination thereof) vibrational spectroscopy for IR spectra prediction. MACE-OFF23, in particular, provides excellent performance for molecular geometries and frequencies, making it a promising method for calculating thermostatistical contributions for large systems besides the spectroscopy purposes. For very large systems, exceeding roughly 700 atoms, we expect MACE-OFF23 to gain a computational cost advantage over the xTB methods due to the MLPs better scaling behaviour with system size. Future work should focus on investigating more accurate reference frequencies both by going beyond the B3LYP-3c level of theory, and extending beyond the harmonic approximation to include vibrational anharmonicity. Investigating IR spectra for condensed-phase and large amorphous systems, such as proteins, using MACE-OFF23, is another promising avenue of research. New MLP developments, like the inclusion of proper electrostatics66 will be beneficial for this purpose and could allow IR intensity prediction without the need for separate dipole models. Additionally, exploring other spectroscopic techniques, such as Raman spectroscopy by examining polarizability instead of dipole moments, will be valuable.

### v2

@@ 3.4 Performance for simulation of gas-phase IR spectra
For the IR7193 set, we investigated all combinations of the discussed levels of theory. Two sets of reference data were employed: unscaled B3LYP-3c and the respective mass-scaled results, which closely correspond to experimental data.21 For brevity, the following discussion focuses on the $r_{\mathrm{msc}}$ matchscore metric, with other metrics provided in the Supporting Information. Results are summarized in Figures 8 to 10, illustrating the average $r_{\mathrm{msc}}$ over IR7193 for all method combinations, referencing different frequency scaling for either reference or actual data.
@@
Figure 8 presents the average $r_{\mathrm{msc}}$ with respect to unscaled B3LYP-3c reference spectra. As expected, there is a correlation between previously observed performance for frequencies, dipole moment prediction, and the similarity scores for IR spectra. All MACE-OFF23 variants consistently outperform the GFNn-xTB methods, regardless of the origin of the dipole data employed. The best average $r_{\mathrm{msc}}$, with a value of 0.813, is provided by MACE-OFF23(large) in combination with MACE-$\mu$ based IR intensities. Combinations of MACE-OFF23 with dipole derivatives obtained from either of the xTB methods also show high $r_{\mathrm{msc}}$ values, around 0.79. Unfortunately, IR intensities obtained via the classical EEQ model show much poorer performance, reflecting the previously observed lack of accuracy in dipole moment prediction. The semi-classical CEH model performs between the xTB and EEQ-derived IR predictions, contrary to the similar behavior of this method for dipole moments compared to GFN1-xTB (see section 3.3). A comparable performance, moderately improving on CEH, is observed for IR spectra with intensities predicted by the kQEq method. As observed for the dipole moments above, kQEq significantly improves upon the classical EEQ models, but the underlying methodical limitations prevent reaching the accuracy of a self-consistent semiempirical treatment. However, note again that kEQq was trained only for the HCNOF element subset and matchscore averages are not fully comparable to the other methods. In general, similar trends are observed for the combination of GFNn-xTB frequencies with various dipole moment methods: MACE-$\mu$ and xTB-derived IR intensities are accurate, while classical and CEH-based IR intensities are not. Employing a force-field, even one trained for frequency prediction such as GFN-FF, notably worsens the obtained IR spectra. Multiple factors probably contribute to this performance. Most importantly, GFN-FF has a significant number of large RMSD minima after optimisation, as discussed in section 3.1, which naturally affects the predicted frequencies and dipoles and hence must contribute to the poor results. The best predictions are again obtained in combination with GFNn-xTB or MACE-$\mu$ calculated IR intensities, although the assocuated computational costs exceed those for calculating the GFN-FF Hessian.
Employing mass-scaled reference data to better reproduce experimental gas-phase spectra21 has a significant impact on the observed performance, as shown in Figure 9.
@@
All MACE-OFF23-based calculations show substantially poorer $r_{\mathrm{msc}}$ values, regardless of the employed dipole moment methods, which can be clearly assigned to the shift in frequencies. On the other hand, all GFN methods are much less affected by the changing reference. GFN2-xTB performs better than its predecessor GFN1-xTB, and even has slightly higher $r_{\mathrm{msc}}$ than with the plain B3LYP-3c reference. Apparently, the formulation of the GFN method family is more consistent in reproducing the molecular frequencies (as observed already for the ZPVE, see section 3.2), and hence preform similarly in both cases. This result shows that the GFN methods are a reasonable choice for inexpensively calculating frequencies, since, on average, robust performance is expected. Unfortunately, this observation also implies a general inability of the GFN methods to achieve much better IR spectra predictions, as well as potential shortcomings of the comparison metric, which is discussed below.
In Figure 10 both the reference data and the low-cost simulations target IR spectra closely corresponding to the experiment.21
@@
Compared to Figure 9, the behaviour of MACE-OFF23-based predictions with regard to the mass-scaled B3LYP-3c reference can be rectified by employing the linear frequency scaling factors obtained in section 3.2. In combination with either the xTB-based dipole derivatives or MACE-$\mu$, all MACE-OFF23 MLPs consistently achieve average $r_{\mathrm{msc}}$ over 0.8, with the overall best benchmark of 0.859 for MACE-OFF23(large)+MACE-$\mu$. A similar trend is seen for GFN1-xTB and GFN2-xTB, with the latter theory performing better. The best average $r_{\mathrm{msc}}$ among these SQM methods is 0.787, which is achieved with GFN2-xTB in combination with MACE-$\mu$. Combining the MACE-OFF23 frequencies with kQEq-based IR intensities generally passes a $r_{\mathrm{msc}}$ of 0.71, which makes them roughly comparable to GFN1-xTB IR spectra employing GFN2-xTB dipole derivatives, moderately improving upon the CEH results. As before, GFN-FF-based predictions of IR spectra, as well as all IR spectra calculated with underlying EEQ or CEH dipole moment derivatives, are lacking accuracy. In fact, GFN-FF results scaled by the linear factor determined via the ZPVE are on average worse than both these alternatives, as seen in Figure 8 and Figure 9, which is further testimony to an asystematic behaviour of the method. For the underlying dipole moment derivatives, classical charge-equilibrium models exhibit relatively poor average $r_{\mathrm{msc}}$ values throughout, independent of both reference values and methods used for frequency calculation. Hence, their use in IR spectra computation should generally be avoided. The semiclassical CEH model, on the other hand, while certainly worse than either a SQM or specially trained MLP, shows promising performance. Results for this method could certainly be improved with a specialized parametrization, and could motivate further development of similar single-diagonalization SQM Hamiltonians.
@@
Unfortunately, these results can also be interpreted as a lack of precision in matchscore-based metrics like $r_{\mathrm{msc}}$. While clearly defined or linearly shifted spectra are sufficiently well compared by matchscore metrics, for asystematically behaving methods the comparison is too insensitive to accurately distinguish spectra, and will produce some average $r_{\mathrm{msc}}$ value (as seen for GFN1-xTB and GFN2-xTB, as well as their combination with MACE-$\mu$ in Figure 9). A better metric should include clear assignment of vibrational modes between reference and calculated spectra, for which, to the best of our knowledge, no automated procedure exists, although attempts exist to project the IR intensity along localized modes and use an appropriate assignment accordingly.66 Furthermore, such a metric would only work if it is known a priori that the same conformation of the same molecule is to be compared and information about the vibrational modes is accessible for the reference spectra. If the goal is, for example, the identification of unknown compounds from experimental measurements, only matchscore metrics seem suitable. Hence, due to lack of better alternatives, we continue to refer to $r_{\mathrm{msc}}$, but advise caution to the reader. On the other hand, obtaining generally mediocre values of $r_{\mathrm{msc}}$ as seen for GFNn-xTB, may signify suitability in applications where an exact order of the frequencies and modes is not overly important, for example the calculation of the ZPVE or vibrational entropy. This fitness is in line with the original GFNn-xTB design purpose13 and hence seems to be confirmed by our results. For a visual assessment of $r_{\mathrm{msc}}$, examples of a “good”, “average” and “mediocre” matchscore are given in Figure 11, depicted as an overlay between the mass-scaled B3LYP-3c reference and the MACE-OFF23(large)+MACE-$\mu$ computed IR spectra with applied ZPVE scaling factor.
@@
In general, matchscores of 0.9 or higher exhibit a close match between the two compared spectra, with only minor differences in relative intensities. Scores between 0.8 and 0.9 likewise denote good correspondence of spectra, although some frequency shifted signals may appear, as seen for modes around 3000 to 3500 cm-1 in Figure.11b. Spectra with $r_{\mathrm{msc}}$ below 0.8 commonly disagree with regard to at least one signal, as for the OH vibrational mode at around 3500 cm-1 in Figure 11c, where both the frequency and intensity are overestimated by the low-cost computation. A mismatch of this type is a common cause of “mediocre” or worse matchscore results, despite other regions of the spectrum (e.g. the fingerprint regime up to $\approx$1500 cm-1) exhibiting good agreement. Spectra with overall fewer IR signals are more prone to this effect, while signals in cluttered spectra will often overlap regardless, and hence contribute to $r_{\mathrm{msc}}$. For benchmark purposes with theory-to-theory comparisons being made, a new metric depending on identification and matching of the actual vibrational mode would clearly be preferable.
@@ 4 Conclusion
In this work, we tested combinations of low-cost semiempirical, force-field, and MLP methods for calculating gas-phase IR spectra within the doubly-harmonic approximation. Our approach involves benchmarking the GFN1-xTB, GFN2-xTB, GFN-FF and MACE-OFF23 models for calculating molecular geometries and associated harmonic frequencies, as well as the GFN1-xTB, GFN2-xTB, CEH, EEQ and MACE-$\mu$ models for molecular dipole moments. Our analysis focuses mostly on the IR7193 set, collecting some 7193 organic molecules and their IR spectra. Linear frequency scaling factors were determined based on the ZPVE fraction relative to a specially (mass-)scaled theoretical reference, B3LYP-3c. IR spectra were obtained by combining frequency predictions with IR intensities calculated via numerically derived dipole moment derivatives.
@@
A key observation is that the MACE-OFF23 MLPs are found to be particularly effective for drug-like molecules, providing accurate geometries, frequencies, and IR spectra that outperform the GFNn-xTB methods. Interestingly, MACE-OFF23 appears to inherit characteristics similar to DFT methods, as evidenced by the comparable importance and magnitude of harmonic frequency scaling factors. This result is contrasts with the evaluated semiempirical methods, which are mostly asystematic with regard to such treatments, and on average are only little affected by a linear scaling of the harmonic frequencies. In general, although the GFNn-xTB methods offer decent performance and speed for smaller system sizes and inherently include dipole moments, they often do not match the accuracy of MACE-OFF23. This result is clearly seen from the best observed average $r_{\mathrm{msc}}$ of 0.787 of GFN2-xTB versus a $r_{\mathrm{msc}}$ of 0.859 for MACE-OFF23(large), both in combination with MACE-$\mu$-based IR intensities. Both MACE-$\mu$ and GFNn-xTB exhibit reasonable performance for calculation of the underlying dipole moment derivatives, with the SQM methods having a computational cost advantage for drug-sized molecules. The use of classical EEQ models for this purpose should be avoided. While CEH does not reach the accuracy of SQM or MLP dipole predictions, the observed performance is a clear motivator for future research.
@@
In summary, we see an opportunity for both MLP and SQM-based prediction of IR vibrational spectra, or a combination of techniques. MACE-OFF23, in particular, provides excellent performance for molecular geometries and frequencies, making it a promising method for calculating thermostatistical contributions for large systems in addition to the spectroscopic applications. For very large systems, exceeding roughly 700 atoms, we expect MACE-OFF23 to gain a computational cost advantage over the xTB methods due to the MLPs better scaling behaviour with system size. Future work should focus on investigating more accurate reference frequencies both by going beyond the B3LYP-3c level of theory, and extending beyond the harmonic approximation to include vibrational anharmonicity. Investigating IR spectra for condensed-phase and large amorphous systems, such as proteins, using MACE-OFF23, is another promising avenue of research. New MLP developments, especially the inclusion of proper electrostatics,67 will be beneficial for this purpose and could allow IR intensity prediction without the need for separate dipole models. Additionally, exploring other spectroscopic techniques, such as Raman spectroscopy by examining polarizability instead of dipole moments, will be valuable.
Glossary
Abbreviation Meaning
BSSE Basis set superposition error
B3LYP-3c short notation for B3LYP-D3(BJ)ATM-gCP/def2-mSVP
CEH Charge extended Hückel model
CN Coordination number
CREST Conformer-Rotamer Sampling Tool
DFT Density functional theory
DHA Doubly-harmonic approximation
EEQ Charge equilibrium model
FF Force-field
GFN Geometries, Frequencies, Non-covalent interactions
GFN-FF GFN force-field
IR Infrared (spectroscopy)
kQEq Kernel charge equilibration model
MACE Higher Order Message Passing Atomic Cluster Expansion
MACE-OFF23 MACE organic force-field
MACE-$\mu$ MACE dipole model
MAD Mean absolute deviation
MAE Mean absolute error
MD Mean deviation
ML Machine learning
MLP Machine learning potential
PEL Potential energy landscape (synonym to PES)
PES Potential energy surface (synonym to PEL)
RMSD Root-mean-square-deviation (of atomic positions)
RMSE Root-mean-square-error
SPICE Small-molecule/Protein Interaction Chemical Energies dataset
SQM Semiempirical quantum mechanics
xTB eXtended Tight-Binding
ZPVE Zero-point vibrational energy